# Counterspeech the ultimate shield! Multi-Conditioned Counterspeech Generation through Attributed Prefix Learning

### v2

@@ 5.1 Baselines
To evaluate the efficacy of various models, we investigate Generate Prune Select (GPS) Zhu and Bhat (2021), a three-stage pipeline encompassing autoencoding, grammatical filtering, and response selection. Furthermore, we optimize DialoGPT Zhang et al. (2020b) to produce contextually coherent responses. Additionally, we incorporate CoARL Hengle et al. (2024), the state-of-the-art intent-conditioned CS generation approach. Furthermore, we explore prefix tuning on Vanilla FLAN-T5XXL, and BARTLarge (excluding HiPPrO) and conduct experiments with HIPO (without preference optimization) using different virtual token (VT) sizes (VT = $3,5,7,10$), VT = $3$ emerging as the optimal configuration in terms of both parameter efficiency ($589,824$ trainable parameters, $0.0052\%$ of total model parameters) and performance metrics. We consider VT = $3$ for preference tuning. With same setup, we experiment with decoder-only models like GPT2XL, Llama 3.1 Instruct8B and Mixtral Instruct7B. We also utilize DPO to enhance our evaluation further. Our comprehensive assessment encompasses both zero-shot and few-shot performances on three LLMs: Vanilla FLAN-T5XXL Chung et al. (2022), GPT-3.5-Turbo (ChatGPT) and GPT-4 Ouyang et al. (2022b) (see Appendix 8.6 and 8.7). In all the above setups we do not consider natural language instruction prompts.
@@ 5.2 Evaluation Metrics
Evaluating CS generation presents challenges due to its dynamic nature, diverse response possibilities, and absence of standardized metrics Chung et al. (2023), prompting our framework to utilize comprehensive, multi-dimensional evaluation metrics. These evaluation metrics include lexical similarity, semantic similarity or relevance, intent conformity, emotion conformity, target conformity, and toxicity score. Lexical similarity is evaluated using Rouge Lin (2004) and Meteor (M) Banerjee and Lavie (2005), which quantify the linguistic alignment between generated and reference texts. Semantic relevance is assessed with cosine similarity (CoSim) Reimers and Gurevych (2019) and BERTScore (BS) Zhang et al. (2020a), ensuring that generated CS engages meaningfully with the primary topic of hate speech. A low relevance score implies a lack of topical coherence, where the CS fails to adequately address the primary subject of hate speech. We also evaluate the effectiveness of incorporating intention, emotional resonance, and target alignment through Intent Conformity (IC) Gupta et al. (2023), Emotion Conformity (EC), and Target Conformity (TC), respectively. These metrics are particularly valuable in scenarios where ground-truth counterspeech is unavailable for unseen hate speech instances. We first train three distinct RoBERTa-large models on our dataset to measure the IC, EC, and TC scores. The models achieve testing accuracy of $0.86$, $0.75$, and $0.88$, respectively, before being considered as evaluation metrics. Toxicity (T) [^6] https://www.perspectiveapi.com/ levels of generated CS using the Hanu and Unitary team (2020) library, ensure that our approach promotes respectful and safe communication.
@@ 6.1 Quantitative Results
The generated CS should ideally be a balanced generation of both intent and emotional attributes, ensuring that responses not only address the harmful content effectively but also align with intent and emotional tone along with the target audience. However, the evaluation results indicate that it is a challenge for many models to achieve this balance. For instance, while GPT-3.5-Turbo ZS demonstrates a low IC score of $0.533$, its EC score is significantly high at $0.783$, highlighting the difficulty in generating CS that is both intent-aligned and emotionally resonant. Similarly, GPT-4 ZS shows a high EC score of $0.790$, yet its IC score of $0.545$ suggests that even more advanced LLMs struggle to maintain emotional alignment in the zero-shot and few-shot scenarios. In contrast, HiPPrO demonstrates a better balance, achieving an IC score of $0.929$ and an EC score of $0.706$. HiPPrO’s IC score is substantially higher, indicating its superior ability to generate intent-aligned CS while still maintaining a better level of emotional resonance. This suggests that HiPPrO is more adept at maintaining the stability between these two critical attributes during the generation process.
@@ 6.2 Ablation Study
Our ablation studies also include evaluations of Vanilla FLAN-T5XXL and Vanilla BARTLarge, both with prefix tuning, without hierarchical learning, providing further insights into HiPPrO’s effectiveness. We observe that while these models offer competitive performance, they do not surpass HiPPrO in several key metrics. For Vanilla FLAN-T5XXL, prefix tuning yields ROUGE-1, ROUGE-2, and ROUGE-L scores of $0.229$, $0.052$, and $0.158$, respectively. These scores are slightly lower than those of HiPPrO, indicating that while the prefix-tuned FLAN-T5XXL performs well, HiPPrO’s output is more detailed and comprehensive. However, the IC scores for Vanilla FLAN-T5XXL and BARTLarge are $0.470$ and $0.249$, respectively, while their EC scores are $0.666$ and $0.453$, respectively. HiPPrO achieves higher IC and EC scores ($0.929$ and $0.706$ respectively), highlighting the limitations of simple prefix-tuning methods in effectively aligning CS with multiple attributes.
@@ 6.3 Human Evaluation
Previous studies (Jones et al., 2024; Wang et al., 2023) emphasize the need for a dual evaluation framework, as automatic metrics show weak correlation with human judgments of counterspeech effectiveness. A comprehensive human evaluation was conducted on a random subset of responses from the top-performing CS generation methods (HiPPrO, CoARL, Few-Shot GPT-4, and GPT-3.5 Turbo), ensuring uniform distribution across intents and emotions. A diverse panel of 35 experts in NLP and social sciences (aged 20-35, 45% male, 55% female) evaluated and ranked these responses based on several key metrics. We followed Hengle et al. (2024) for our human evaluation. The evaluation framework consists of five key metrics: Independent Counterspeech (ICS) to gauge the response’s self-sufficiency; Adequacy (Ad) to assess its linguistic quality; Contextual Relevance (CoRl) to measure its responsiveness to hate speech components; and Argumentative Effectiveness (ArgE) to evaluate its carefulness and convincing. We present the comparative performance of HiPPrO against leading methods through Win Rate scores (see Table 3). For ICS, HiPPrO outperforms the baselines with win rates of $0.91$ and $0.89$ over GPT-4 and GPT-3.5, showing its ability to generate CS that operates effectively without needing extra context. Regarding Ad, HiPPrO achieves higher scores of $0.85$ and $0.87$, reflecting the superior grammatical accuracy and fluency of CS. In terms of CoRl, HiPPrO’s win rates of $0.89$ and $0.87$ highlight its strength in addressing crucial aspects of hate speech, such as targeted biases. Finally, for ArgE, HiPPrO leads with scores of $0.89$ and $0.90$, indicating its effectiveness in delivering compelling and well-structured CS. These results collectively show HiPPrO’s robust performance across all metrics compared to the baseline models.
@@ 7 Conclusion
This study presented HiPPrO, a novel two-stage framework for generating controllable, multi-attributed counterspeech. The initial stage comprised a hierarchical learning process, where the model acquired attribute-specific prefixes, thereby guiding the LLM towards targeted counterspeech generation. The subsequent stage involved refining the outputs to enhance their human-like quality and non-toxicity, employing a reward- and reference-free alignment approach. Additionally, we introduced the multiCONAN dataset with intent- and emotion-specific counterspeech. An extensive evaluation, incorporating a range of quantitative and qualitative measures, demonstrated HiPPrO’s superiority over multiple baselines.
@@
Our research presents several limitations that warrant consideration. Firstly, the dataset utilized for hate speech and counterspeech is not comprehensive, potentially omitting various forms and targets of online hate. Secondly, the framework’s reliance on pre-trained models may introduce inherent biases or inaccuracies stemming from these source models. Additionally, the evaluation metrics employed do not fully align with human perceptions of counterspeech quality, thereby failing to capture the intricate nuances of natural language interactions. Moreover, our framework does not address the possibility of feedback loops or escalation that could arise following the generation of counterspeech, which may influence the long-term effectiveness and impact of our approach. Lastly, while efforts were made to maintain high-quality annotations for counterspeech, it is conceivable that our dataset may not match the caliber of those annotated by more experienced operators from NGOs, such as those found in the Multi-Target CONAN Fanton et al. (2021b) dataset. Future research could mitigate these limitations by expanding and diversifying the dataset, enhancing the reward function and evaluation criteria, and incorporating dialogue modeling into the framework.

### v3

@@ 5.1 Baselines
To evaluate the efficacy of various models, we investigate Generate Prune Select (GPS) Zhu and Bhat (2021), a three-stage pipeline encompassing autoencoding, grammatical filtering, and response selection. Furthermore, we optimize DialoGPT Zhang et al. (2020b) to produce contextually coherent responses. Additionally, we incorporate CoARL Hengle et al. (2024), the state-of-the-art strategy-conditioned CS generation approach. Furthermore, we explore prefix tuning on Vanilla FLAN-T5XXL, and BARTLarge (excluding HiPPrO) and conduct experiments with HIPO (without preference optimization) using different virtual token (VT) sizes (VT = $3,5,7,10$), VT = $3$ emerging as the optimal configuration in terms of both parameter efficiency ($589,824$ trainable parameters, $0.0052\%$ of total model parameters) and performance metrics. We consider VT = $3$ for preference tuning. With same setup, we experiment with decoder-only models like GPT2XL, Llama 3.1 Instruct8B Grattafiori et al. (2024), Mixtral Instruct7B Jiang et al. (2024), DeepSeek-R1-Distill-Llama-8B and DeepSeek-llm-7b-chat DeepSeek-AI et al. (2025). We also utilize DPO to enhance our evaluation further. Our comprehensive assessment encompasses zero-shot and few-shot performances on three LLMs: Vanilla FLAN-T5XXL Chung et al. (2022), GPT-3.5-Turbo (ChatGPT), and GPT-4 Ouyang et al. (2022b) (see Appendix 9.6 and 9.7) and simple retrieval-based methods using state-of-the-art open-source LLMs, employing ‘faiss’Douze et al. (2025) as the retrieval method to retrieve the top five training counterspeech examples (see Appendix 9.8).
@@ 5.2 Evaluation Metrics
Evaluating CS generation presents challenges due to its dynamic nature, diverse response possibilities, and absence of standardized metrics Chung et al. (2023), prompting our framework to utilize comprehensive, multi-dimensional evaluation metrics. These evaluation metrics include lexical similarity, semantic similarity or relevance, strategy conformity, emotion conformity, target conformity, and toxicity score. Lexical similarity is evaluated using Rouge Lin (2004) and Meteor (M) Banerjee and Lavie (2005), which quantify the linguistic alignment between generated and reference texts. Semantic relevance is assessed with cosine similarity (CoSim) Reimers and Gurevych (2019) and BERTScore (BS) Zhang et al. (2020a), ensuring that generated CS engages meaningfully with the primary topic of hate speech. A low relevance score implies a lack of topical coherence, where the CS fails to adequately address the primary subject of hate speech. We also evaluate the effectiveness of incorporating strategic, emotional resonance, and target alignment through Strategy Conformity (SC) Gupta et al. (2023), Emotion Conformity (EC), and Target Conformity (TC), respectively. These metrics are particularly valuable in scenarios where ground-truth counterspeech is unavailable for unseen hate speech instances. We first train three distinct RoBERTa-large models on our dataset to measure the SC, EC, and TC scores. The models achieve testing accuracy of $0.86$, $0.75$, and $0.88$, respectively, before being considered as evaluation metrics. Toxicity (T) [^3] https://www.perspectiveapi.com/ levels of generated CS using the Hanu and Unitary team (2020) library, ensure that our approach promotes respectful and safe communication.
@@ 6.1 Quantitative Results
The generated CS should ideally be a balanced generation of both strategy and emotional attributes, ensuring that responses not only address the harmful content effectively but also align with strategy and emotional tone along with the target audience. However, the evaluation results indicate that it is a challenge for many models to achieve this balance. For instance, while GPT-3.5-Turbo ZS demonstrates a low SC score of $0.533$, its EC score is significantly high at $0.783$, highlighting the difficulty in generating CS that is both strategy-aligned and emotionally resonant. Similarly, GPT-4 ZS shows a high EC score of $0.790$, yet its SC score of $0.545$ suggests that even more advanced LLMs struggle to maintain emotional alignment in the zero-shot and few-shot scenarios. In contrast, HiPPrO demonstrates a better balance, achieving an SC score of $0.929$ and an EC score of $0.706$. HiPPrO’s SC score is substantially higher, indicating its superior ability to generate strategy-aligned CS while still maintaining a better level of emotional resonance. This suggests that HiPPrO is more adept at maintaining the stability between these two critical attributes during the generation process.
@@ 6.2 Ablation Study
Our ablation studies also include evaluations of Vanilla FLAN-T5XXL and Vanilla BARTLarge, both with prefix tuning, without hierarchical learning, providing further insights into HiPPrO’s effectiveness. We observe that while these models offer competitive performance, they do not surpass HiPPrO in several key metrics. For Vanilla FLAN-T5XXL, prefix tuning yields ROUGE-1, ROUGE-2, and ROUGE-L scores of $0.229$, $0.052$, and $0.158$, respectively. These scores are slightly lower than those of HiPPrO, indicating that while the prefix-tuned FLAN-T5XXL performs well, HiPPrO’s output is more detailed and comprehensive. However, the SC scores for Vanilla FLAN-T5XXL and BARTLarge are $0.470$ and $0.249$, respectively, while their EC scores are $0.666$ and $0.453$, respectively. HiPPrO achieves higher SC and EC scores ($0.929$ and $0.706$ respectively), highlighting the limitations of simple prefix-tuning methods in effectively aligning CS with multiple attributes.
@@ 6.3 Human Evaluation
Previous studies (Jones et al., 2024; Wang et al., 2023; Hengle et al., 2025) emphasize the need for a dual evaluation framework, as automatic metrics show weak correlation with human judgments of counterspeech effectiveness. A comprehensive human evaluation was conducted on a random subset of $30$ responses from the top-performing CS generation methods (HiPPrO, CoARL, Few-Shot GPT-4, and GPT-3.5 Turbo) with a random seed value of $1$, ensuring uniform distribution across strategies and emotions. A diverse panel of 35 experts in NLP and social sciences (aged 20-35, 45% male, 55% female) evaluated and ranked these responses based on several key metrics. We followed Hengle et al. (2024) for our human evaluation. The evaluation framework consists of five key metrics: Independent Counterspeech (ICS) to gauge the response’s self-sufficiency; Adequacy (Ad) to assess its linguistic quality; Contextual Relevance (CoRl) to measure its responsiveness to hate speech components; and Argumentative Effectiveness (ArgE) to evaluate its carefulness and convincing. We present the comparative performance of HiPPrO against leading methods through Win Rate scores (see Table 3). For ICS, HiPPrO outperforms the baselines with win rates of $0.91$ and $0.89$ over GPT-4 and GPT-3.5, showing its ability to generate CS that operates effectively without needing extra context. Regarding Ad, HiPPrO achieves higher scores of $0.85$ and $0.87$, reflecting the superior grammatical accuracy and fluency of CS. In terms of CoRl, HiPPrO’s win rates of $0.89$ and $0.87$ highlight its strength in addressing crucial aspects of hate speech, such as targeted biases. Finally, for ArgE, HiPPrO leads with scores of $0.89$ and $0.90$, indicating its effectiveness in delivering compelling and well-structured CS. These results collectively show HiPPrO’s robust performance across all metrics compared to the baseline models.
@@ 7 Conclusion
This study presented HiPPrO, a novel two-stage framework for generating controllable, multi-attributed counterspeech. The initial stage comprised a hierarchical learning process, where the model acquired attribute-specific prefixes, thereby guiding the LLM towards targeted counterspeech generation. The subsequent stage involved refining the outputs to enhance their human-like quality and non-toxicity, employing a reward- and reference-free alignment approach. Additionally, we introduced the multiCONAN dataset with strategy- and emotion-specific counterspeech. An extensive evaluation, incorporating a range of quantitative and qualitative measures, demonstrated HiPPrO’s superiority over multiple baselines.
@@
Our research presents several limitations that warrant consideration. Firstly, the dataset utilized for hate speech and counterspeech is not comprehensive, potentially omitting various forms and targets of online hate. Secondly, the framework’s reliance on pre-trained models may introduce inherent biases or inaccuracies stemming from these source models. Additionally, the evaluation metrics employed do not fully align with human perceptions of counterspeech quality, thereby failing to capture the intricate nuances of natural language interactions. Moreover, our framework does not address the possibility of feedback loops or escalation that could arise following the generation of counterspeech, which may influence the long-term effectiveness and impact of our approach. Lastly, while efforts were made to maintain high-quality annotations for counterspeech, it is conceivable that our dataset may not match the caliber of those annotated by more experienced operators from NGOs, such as those found in the Multi-Target CONAN Fanton et al. (2021b) dataset. Future research could mitigate these limitations by expanding and diversifying the dataset, enhancing the evaluation criteria, and incorporating dialogue modeling into the framework. Our primary focus was on generating effective and non-toxic counterspeech, following prior work Hengle et al. (2024). While we did not explicitly analyze potential biases in this study, we acknowledge that large language models trained on social media data can amplify biases. Addressing and mitigating such biases is indeed a critical area of research.
@@
8 Acknowledgments
We extend our gratitude to the central HPC facility (Padum) at IIT Delhi for computing. We also sincerely thank Logically and Anusandhan National Research Foundation (DST-NSF Grant: DST/INT/USA/NSF-DST/Tanmoy/P-2/2024) for financial support. Tanmoy acknowledges the support of Rajiv Khemani Young Faculty Chair Professorship in Artificial Intelligence.